OpenStudy (anonymous):

Which placement of grouping symbols in the expression gives it a value of 14? 6 • 4 – 4 + 9 – 3 A. 6 • (4 – 4 + 9) – 3 B. 6 • [(4 – 4 + 9) – 3] C. 6 • 4 – [4 + (9 – 3)] D. (6 • 4 – 4) + (9 – 3)

OpenStudy (anonymous):

help me plz

OpenStudy (anonymous):

thinkig

OpenStudy (anonymous):

c

OpenStudy (anonymous):

need explnation?

OpenStudy (anonymous):

Sorry answer is C.6x4-[4+(9-3)]

OpenStudy (anonymous):

ty :D

OpenStudy (anonymous):

can u show me how u got it ?

OpenStudy (anonymous):

@ddDemon_666

OpenStudy (anonymous):

sorry yeah

OpenStudy (anonymous):

6x4=24 9-3=6+4=10 24-10=14 6x4-[4+(9-3] Parenthesis first so 9-3 is 6. then brackets, what is left in the brackets? well [4+6] which is 10, then whats left reading from left to right its 6x4-10 which is also 24-10=14, helped?

OpenStudy (anonymous):

ty :D

OpenStudy (anonymous):

can u help me with another one?

OpenStudy (anonymous):

Ill try

OpenStudy (anonymous):

What is the value of the expression? 40 ÷ [24 – 4 • (2 + 3)]

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

What is the value of the expression? (5 + 10) ÷ 5

OpenStudy (anonymous):

this one

OpenStudy (anonymous):

mmhm.

OpenStudy (anonymous):

3

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Sorry I have a question what do the "?" in side the diamonds represent?

OpenStudy (anonymous):

i did that one

OpenStudy (anonymous):

What is the value of the expression? (5 + 10) ÷ 5

OpenStudy (anonymous):

I dont recognize?

OpenStudy (anonymous):

(5+10) /5

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

THE / is divide

OpenStudy (anonymous):

Yeah i got that after um the answer is 3

OpenStudy (anonymous):

What is the value of the expression? 40 ÷ [24 – 4 • (2 + 3)]

OpenStudy (anonymous):

So those boxes are "/"?

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

they are

OpenStudy (anonymous):

40/[24-4x(2+3)]

OpenStudy (anonymous):

the / is divide

OpenStudy (anonymous):

sorry is the 4x "4x" or is the 4 multiplying?

OpenStudy (anonymous):

Hello?